Given that the complex number z = -2 + 7i is a root to the equation:

z³ + 6 z² + 61 z + 106 = 0


find the real root to the equation.


1
Expert's answer
2021-03-16T07:16:49-0400

Since the complex number "z = -2 + 7i" is a root of the equation "z^3 + 6 z^2 + 61 z + 106 = 0" with real coefficients, the conjugate complex number "\overline{z} = -2 - 7i" is also a root of the equation. By

Bézout's theorem, "(z+2-7i)(z+2+7i)=(z+2)^2+49=z^2+4z+53" divides "z^3 + 6 z^2 + 61 z + 106". It follows that "z^3 + 6 z^2 + 61 z + 106=(z^2+4z+53)(z+2)". We conclude that "z+2=0", and hence "z=-2" is a real root of the equation "z^3 + 6 z^2 + 61 z + 106 = 0."